# VideoGen-of-Thought: A Collaborative Framework for Multi-Shot Video Generation

### v1

@@ 5.5 Ablation Studies
In this section, we present ablation studies to evaluate the impact of two key components: (1) Enhanced Prompt (EP) Generation, and (2) Identity-Preserving (IP) Control for keyframe consistency. Both quantitative and qualitative comparisons of VGoT are conducted by systematically removing either EP or IP modules, or both, using the user input: A set of one-sentence prompts, 30 shots, describing the journey of Carlos, from a young boy on his first bike to becoming a celebrated world champion cyclist.
The qualitative results, shown in Figure 5, illustrate that the complete VGoT framework (top row) maintains both character identity and narrative depth, while systematically removing EP (second row), IP (third row), or both (bottom row) results in a noticeable decline in visual diversity, narrative coherence, and character consistency.
The quantitative results, as presented in Table 3, further support our visual findings, with the highest PSNR and IS values indicating superior quality for our complete model. Two key insights from Table 3 are worth noting:
1. The full model achieves the lowest CLIP Score across all ablations. This is because the EP module enriches textual descriptions and the IP module preserves consistent character portrayal, resulting in generated visuals that deviate from the original prompts but ultimately enhance the video quality.
2. The removal of EP leads to improvements in FC and SC scores. However, while these metrics reflect increased cross-shot consistency, they do not capture the overall logic and diversity of the storyline. Comparing the middle and right portions of the top and second rows in Figure 5 shows that the absence of EP results in diminished narrative diversity, with visually similar shots benefiting consistency metrics but compromising the story’s richness. These ablation studies clearly demonstrate the importance of each component in producing high-quality, logically sound, and consistent multi-shot videos.
Figure 5: Visual Demonstration of the ablation studies of VGoT
@@ 6 Conclusion
In this paper, we introduce VideoGen-of-Thought (VGoT), a modular framework that tackles the challenges of generating multi-shot videos with narrative coherence, visual consistency, and stylistic fidelity. VGoT breaks down video generation into four collaborative modules, ensuring high semantic alignment, keyframe consistency, smooth transitions, and scene-specific guidance. Our evaluations demonstrate that VGoT outperforms existing baselines in maintaining flow consistency, face coherence, and stylistic alignment, offering a promising approach for scalable and coherent multi-shot video generation.
@@
Appendix A Implementation Details of Script Module
@@
The Script Module plays a fundamental role in converting high-level user input into a detailed series of prompts for each shot within the multi-shot video generation process. The specific process is to convert a single sentence user input $S$, into a more detailed and structured description $S^{\prime}$, which is then decomposed into a set of prompts $P=\{p_{1},p_{2},\dots,p_{N}\}$, corresponding to each of the $N$ shots required for the complete video. This process uniformly adopts a large language model (LLM) and uses prompt engineering to ensure the reasonableness of the generated video. This process enables the generation of a logical, stepwise narrative structure that serves as the backbone for subsequent keyframe and shot-level generation.
@@
For example, consider the user input: e.g., a set of one-sentence prompts, 30 shots, describe a story about three avatars (two friend Mike and Jane, one bad guy Tom) and their fantasy adventure in ancient pyramid.. The LLM would first transform this input into a detailed version $S^{\prime}$ consisting of 30 one-sentence scripts like $s_{1}$: “Mike’s Discovery: Mike examines an ancient map with intense focus, revealing the path to the pyramid.” Subsequently, each $s_{i}$ in $S^{\prime}$ is decomposed into prompts $P$, such as $p_{1}$: Character: Mike, holding an ancient map with Jane by his side. Background: A dense jungle filled with mist and towering trees. Relation: Mike studies the map closely, pointing to a pyramid. Camera Pose: Medium shot focusing on Mike and Jane. HDR Description: Soft light filters through the trees, creating dynamic shadows on the map and characters.”. This structured approach ensures narrative coherence across multiple shots, laying the foundation for the generation process.
Appendix B Additional Results
To further illustrate the effectiveness of our proposed VideoGen-of-Thought (VGoT) framework, we present additional qualitative results showcasing the generated videos. Besides the pyramid story with three avatars provided in Sec A and materials, VGoT achieves a better reasonable storyline and visual consistency in stories about one protagonist’s growth at different times. We prepared ten showcases in the materials, and a large visual representation of VGoT’s output is provided in Figure 6, spanning two full pages to ensure sufficient detail and clarity for each generated shot. This figure demonstrates the stepwise progression of narrative shots, showcasing the seamless visual consistency and logical coherence in a wide variety of generated multi-shot videos.
@@ Appendix B Additional Results
Appendix C User Study
Appendix D Limitations and Future Work
While our VGoT framework offers significant advancements in multi-shot video generation, there are still limitations that need to be addressed, along with directions for future work.
1. Single Identity-Preserving Embedding per Shot: Currently, each shot only receives one identity-preserving (IP) embedding, which limits the flexibility to depict complex scenarios involving multiple characters or subjects. Future work will explore extending the framework to handle multi-subject driven shot-level video generation, thereby enhancing narrative richness and visual diversity.
2. Limited Evaluation Metrics: Our current evaluation focuses mainly on face and style consistency, which may not capture all aspects of multi-shot video quality comprehensively. There is a lack of detailed multi-shot evaluation metrics to assess narrative flow and cross-shot coherence. Future work will involve researching advanced cross-shot metrics and using these metrics to inform and guide model training, potentially improving overall video generation quality.

### v2

@@ 5.3 Ablation Studies
We analyze two core components through systematic removal: (1) Dynamic Storyline Modeling (DSM) and (2) Identity-Preserving Portraits (IPP). Using the 30-shot cycling narrative, Table 3 reveals three critical patterns:
1. CLIP-Logic Tradeoff: The full model achieves lowest CLIP score (0.1111 vs 0.1223 baseline) but highest PSNR (25.79) and IS (7.52), confirming that DSM’s narrative enrichment and IPP’s consistency control prioritize cinematic quality over literal prompt matching.
2. Consistency Costs: Removing DSM boosts $\Omega_{\text{CS-FC}}$ by 20.2% (0.3291 vs 0.2738) and SC by 8.5% (0.4186 vs 0.3859), but as Fig 5 shows, this comes at severe narrative diversity loss—identical camera angles and repetitive scenes dominate DSM-ablated outputs.
3. Component Synergy: IPP alone achieves 0.8745 $\Omega_{\text{WS-FC}}$ (5.3% higher than full model), but combined DSM+IPP delivers optimal balance—8.3% better $\Omega_{\text{CS-SC}}$ than IPP-only versions while maintaining visual quality (25.79 PSNR vs 23.92).
These results prove both components are essential: DSM enables story progression while IPP ensures continuity, together resolving the consistency-diversity paradox in multi-shot generation.
@@ 6 Conclusion
We present VideoGen-of-Thought (VGoT), a structured framework that automates multi-shot video generation with reasonable storylines and visual consistency from one sentence through minimal manual intervention, through three core innovations: dynamic storyline modeling, identity-aware cross-shot propagation, and adjacent latent transition mechanisms. VGoT achieves 20.39% higher within-shot face consistency and 17.36% better style consistency than previous state-of-the-art methods, with over 100% improvements in cross-shot metrics and 10× fewer manual adjustments than alternatives while maintaining director-level narrative flow. Our work redefines automated multi-shot generation, bridging raw visual synthesis with cinematic storytelling through systematic visual stories decomposition.
@@
7 Detailed Example of Dynamic Storylines
@@
Dynamic Storylines Modeling plays a fundamental role in converting high-level user input into a detailed series of prompts for each shot within the multi-shot video generation process. The specific process is to convert a single sentence user input $S$, into a more detailed and structured description $S^{\prime}$, which is then decomposed into a set of prompts $P=\{p_{1},p_{2},\dots,p_{N}\}$, corresponding to each of the $N$ shots required for the complete video. This process uniformly adopts a large language model (LLM) and uses prompt engineering to ensure the reasonableness of the generated video. This process enables the generation of a logical, stepwise narrative structure that serves as the backbone for subsequent keyframe and shot-level generation.
@@
For example, consider the user input: e.g., a set of one-sentence prompts, 30 shots, describe a story about three avatars (two friend Mike and Jane, one bad guy Tom) and their fantasy adventure in an ancient pyramid.. The LLM would first transform this input into a detailed version $S^{\prime}$ consisting of 30 one-sentence scripts like $s_{1}$: “Mike’s Discovery: Mike examines an ancient map with intense focus, revealing the path to the pyramid.” Subsequently, each $s_{i}$ in $S^{\prime}$ is decomposed into prompts $P$, such as $p_{1}$: Character: Mike, holding an ancient map with Jane by his side. Background: A dense jungle filled with mist and towering trees. Relation: Mike studies the map closely, pointing to a pyramid. Camera Pose: Medium shot focusing on Mike and Jane. HDR Description: Soft light filters through the trees, creating dynamic shadows on the map and characters.”. This structured approach ensures narrative coherence across multiple shots, laying the foundation for the generation process.
Our self-validation mechanism ensures cinematic rigor through iterative refinement. Consider a draft shot description: 2. Inside the helicopter, a diverse team of scientists, their faces filled with anticipation and anxiety, pore over maps and equipment
Initial Generation Attempt:
$p_{2}^{(1)}=\begin{cases}p_{\text{cha}}&\text{Team of scientists with mixed ages}\\ p_{b}&\text{Helicopter interior}\\ p_{r}&\text{Studying maps}\\ p_{\text{cam}}&\text{Close-up}\\ p_{h}&\text{(Missing HDR specification)}\end{cases}$
Validation fails with $K(p_{2}^{(1)},\mathcal{C}_{\text{film}})=0$ due to incomplete HDR description. The LLM regenerates with lighting constraints:
Validated Output:
$p_{2}^{*}=\begin{cases}p_{\text{cha}}&\text{Diverse team in expedition gear, anxious expressions}\\ p_{b}&\text{Helicopter over Amazon rainforest}\\ p_{r}&\text{Team engrossed in equipment}\\ p_{\text{cam}}&\text{Wide shot showing interior/exterior}\\ p_{h}&\text{Sunlight streaming through windows, warm glow}\end{cases}$
This satisfies both criteria: $C(p_{2}^{*},p_{1})=0.92>\tau_{c}$, $K(p_{2}^{*},\mathcal{C}_{\text{film}})=1$, ensuring director-level coherence from user input under minimal manual intervention.
8 Additional Results
We validate VideoGen-of-Thought (VGoT) through four narrative archetypes. Type 1: Longitudinal Character Development demonstrates decade-spanning consistency using prompts like “30-shot story of Marco discovering ancient secrets through culinary journeys”, where our framework maintains consistency across aging sequences. Type 2: Multi-Actor Scenes handles complex group dynamics in scenarios like “Abandoned factory transformed into community art center”, preserving relationship continuity between complex stories and multiple characters across shots through identity-aware propagation. Type 3: Non-Human Narratives extends identity preservation to fantastical subjects, as shown in “Mycelium networks and mechanical bees restoring ecosystems”, we explore the creatity of VGoT in marvelous entities. We also evaluate the capability of VGoT to create diverse stories with the same input: “an immigrant’s story of moving to a new country, struggling, and eventually finding success as an entrepreneur.” in Type 4 showcases.